OpenStudy (anonymous):

A square piece of cheese covers 8 square inches of a plate. What is the approximate length of one side of the piece of cheese? (Approximate to the nearest hundredth inch.) 2.81 inches 2.82 inches 2.83 inches 2.84 inches

OpenStudy (bibby):

The area of the square is 8 in^2 What is the formula for square area

OpenStudy (anonymous):

uhhhh

OpenStudy (anonymous):

i accutally don't remember

OpenStudy (bibby):

it's side * side or side^2

OpenStudy (anonymous):

oh

OpenStudy (bibby):

so what do we have? 

OpenStudy (anonymous):

so the length of one side would be 2.82

OpenStudy (bibby):

how'd you get that?

OpenStudy (anonymous):

i was thinking that because accutally im not sure how

OpenStudy (bibby):

you can't just guess. You have the formula for area and algebra at your disposal... A=8 a=x*x

OpenStudy (anonymous):

so like 4 x 4 =8, it has something to do like that right?

OpenStudy (bibby):

4+4=8 4x4=16 \(\huge {x^{2}}=8\) \(\huge \sqrt{x^\cancel{2}}=\sqrt 8\)

OpenStudy (bibby):

for the kiddos googling this \(x=\sqrt8\\x=2.828\implies2.83\)